How many three-digit numbers are divisible by 13?
(A) 7 (B) 67 (C) 69 (D) 76 (E) 77

Answer by chosenpoint(26)

Here's a quick answer off the cuff.

Doing simple multiplication, find the highest product of 13 less than 100 (the lowest possible 3-digit number). In this case, the closest you can get is 13*7=91

Now find the highest product of 13 greater than 999 (the highest possible 3-digit number). In this case it's 13*77=1001

So we can now easily see that the products of 13 that are 3 digit numbers are exclusively given as:

{8, 9, 10...75,76,77} or {8 through 77}

To find the actual amount of products in that range, just subtract 8 from 77

77-8=69

There are 69 3-digit numbers divisible by 13, so the answer is C!!! :)